Determining Manufacturing Costs

Wagner Printing Inc. began printing operations on July 1. Jobs 7-01 and 7-02 were completed during the month, and all costs applicable to them were recorded on the related cost sheets. Jobs 7-03 and 7-04 are still in process at the end of the month, and all applicable costs except factory overhead have been recorded on the related cost sheets. In addition to the materials and labor charged directly to the jobs, $860 of indirect materials and $3,100 of indirect labor were used during the month. The cost sheets, in summary form, for the four jobs during the month are as follows:

Job 7-01 Job 7-02
Direct materials 8,860 Direct materials 4,140
Direct labor 3,400 Direct labor 1,800
Factory overhead 1,870 Factory overhead 990
Total 14,130 Total 6,930
Job 7-03 Job 7-04
Direct materials 12,900 Direct materials 2,550
Direct labor 3,800 Direct labor 500
Factory overhead ? Factory overhead ?

Determine each of the following for July:

a. Direct and indirect materials used.

Direct materials used and transferred to Work in Process $
Indirect materials used and transferred to Factory Overhead $

b. Direct and indirect labor used.

Direct labor used and transferred to Work in Process $
Indirect labor used and transferred to Factory Overhead $

c. Factory overhead applied (a single overhead rate is used based on direct labor cost).
$

d. Cost of completed Jobs 7-01 and 7-02.
$

e. Assume that in addition to indirect materials and indirect labor, factory overhead of $1,500 was incurred during July. Determine the overapplied or underapplied overhead for July.
$  
Answer #1

Answer;-

a. Direct and indirect materials used.

Direct materials used and transferred to Work in Process (8,860+4,140+12,900+2,550) $28,450
Indirect materials used and transferred to Factory Overhead $860

b. Direct and indirect labor used.

Direct labor used and transferred to Work in Process (3,400+1,800+3,800+500) $9500
Indirect labor used and transferred to Factory Overhead $3,100

c. Factory overhead applied (a single overhead rate is used based on direct labor cost).

Overhead rate = 1,870 /3,400 = 55%
$9500*55% = 5,225

d. Cost of completed Jobs 7-01 and 7-02.

Cost of completed jobs = 14,130+6,930 = 21,060

$______________________

if you have any query please ask me in comment box i am here to helps you don't give direct Thumbs down.if you satisfied my work give Thumbs UP

*******THANK YOU*******